Circular No. 7580

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001S IN UGC 5491 T. Puckett, Mountain Town, GA; and A. Sehgal, Woodinville, WA,

report the discovery of an apparent supernova (mag 17.8) on an unfiltered CCD frame (limiting mag 20.1) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Feb. 6.26 UT. SN 2001S is located at α = 10^{h}11^{m}57^{s}.20, δ = +58^{o}52′02′′.9 (equinox 2000.0), which is 7′′.7 west and 10′′.0 north of the center of UGC 5491. The new object was confirmed on CCD frames taken on Feb. 7.01. SN 2001S does not appear on Palomar Sky Survey images taken on 1996 Dec. 20 (limiting mag ∼ 21.0), 1997 May 6 (limiting mag ∼ 21.1) and 1954 Jan. 5 (limiting mag ∼ 19.5).

V445 PUPPIS R. W. Russell, D. K. Lynch, and D. Kim, The Aerospace Corpora-

tion; and M. L. Sitko and S. M. Brafford, University of Cincinnati, report on infrared spectroscopy of V445 Pup on Jan. 31.45 UT using BASS on the Infrared Telescope Facility: “The spectrum (in radiance units W cm^{−2} μm^{−1}) revealed only a smooth, featureless continuum that decreased monotonically with increasing wavelengths (λ) between 3 and 13.6 μm. Its slope is much more shallow than the Rayleigh-Jeans tail of a hot blackbody, close to λ^{−1.5}. The spectrum is consistent with thermal emission from gray-emissivity dust whose temperatures ranged from ∼ 250 K to upwards of 1000 K. Infrared magnitudes were L = 2.63, M = 1.78, N = −0.16 (all ± 0.05). The presence of such a strong infrared excess so early in the object’s evolution suggests that this object is a recurrent nova and has undergone previous outbursts and dust-formation events.”

COMET C/2001 C2 (SOHO) Further to IAUC 7573, D. A. Biesecker reports his measurements of a

fairly bright Kreutz sungrazing comet found by S. Hoenig and X.-m. Zhou on C3 coronagraph images posted at the SOHO website. V magnitude estimates by Biesecker: Feb. 6.654 UT, 7.6; 6.696, 7.5; 6.738, 7.1; 6.779, 6.7; 6.821, 6.7; 6.863, 6.4; 6.904, 6.0; 7.013, 5.7; 7.154, 5.2; 7.221, 4.9; 7.263, 4.5; 7.321, 4.5; 7.488, 4.0. Estimated tail lengths: Feb. 6.904, 0^{o}.1; 7.488, ∼ 0^{o}.2. The reduced positions and orbital elements by B. G. Marsden appear on MPEC 2001-C09.

2001 UT  α_{2000}  δ_{2000}
Feb. 6.346  21^{h}52.^{m}5  −18^{o}07′
2001 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green